OpenStudy (shamil98):

A hot bowl of soup cools according to Newton's law of cooling. Its temperature (in degrees Fahrenheit) at time t is given by the function T(t) = 68 + 144e^-0.04t , where t is expressed in minutes. At what time will the soup reach 200F? How do I solve this?

OpenStudy (charlotte123):

T(t) = 68 + 144e^-0.04t ---> So all of those t's in the equation are expressed in minutes right? Or is it just the one in the beginning? O_o

OpenStudy (shamil98):

It's a function char -_-

OpenStudy (charlotte123):

Well excuse me Mr. I steal yer Peach Juice =.=

OpenStudy (charlotte123):

http://openstudy.com/study#/updates/50995ce0e4b02ec0829cb60e <-- TA DA

OpenStudy (shamil98):

I think I'm supposed to set it to equal 200. but I solved that and got 0.21 which is not an answer choice..

OpenStudy (shamil98):

I want an explanation not an answer.

OpenStudy (charlotte123):

There's an explanation in there ya peanut ლ(́◉◞౪◟◉‵ლ)

OpenStudy (shamil98):

Oh nvm I got it. its 2.175 or 2.18 so 2.18 minutes

OpenStudy (shamil98):

When solving i made an error..

OpenStudy (shamil98):

T(t) = 68 + 144e^-0.04t 200 = 68 + 144e^-0.04t 132/144 = e^-0.04t 0.91666666666 = e^-0.04t ln 0.91666666666 / -0.04 = t 2.17528 = t

OpenStudy (shamil98):

you were of no use.

OpenStudy (shamil98):

huehuehuehue

OpenStudy (charlotte123):

You set T(t) = 200 So your new function would be ~ 200 = 68 + 144e^−0.04t Then ya subtract 68 from both sides ~ Getting 132=144e^−0.04t You are solving for t like any other equation, so you want t on one side and the value on the other side ~ Then ya divide both sides by 144 ~

OpenStudy (charlotte123):

Dude Q_Q

OpenStudy (charlotte123):

You cut me off Q_Q

OpenStudy (charlotte123):

</3

OpenStudy (shamil98):

Yeah, Ik that, apparently your supposed to USE THE EXACT VALUE for 132/144 i put 0.917 i guess rounding was a bad idea o-o

OpenStudy (shamil98):

the exact value is 0.91666666666

OpenStudy (charlotte123):

Peanut head T_T

OpenStudy (shamil98):

#mathlogic

OpenStudy (charlotte123):

You know that I helped ya man ლ(́◉◞౪◟◉‵ლ) Just don't wanna admit it ლ(́◉◞౪◟◉‵ლ)

OpenStudy (charlotte123):

THAT'S RIGHT MEDAL ME ლ(́◉◞౪◟◉‵ლ)

OpenStudy (shamil98):

you better medal me or else I get the belt hun

OpenStudy (shamil98):

I SOLVED IT ALL BY MESELF

OpenStudy (charlotte123):

\(\huge\color{CHOOSE}{\frak{You~Are~No~Match~For~Me~Wizard}}\) ლ(́◉◞౪◟◉‵ლ)

OpenStudy (charlotte123):

The............the belt?! ლ(́◉◞౪◟◉‵ლ)

OpenStudy (charlotte123):

DAM POLAR BEAR ლ(́◉◞౪◟◉‵ლ)

OpenStudy (shamil98):

\[\Huge Good~girl.\]i take my leave.

OpenStudy (charlotte123):

Mother flapper ლ(́◉◞౪◟◉‵ლ)

OpenStudy (charlotte123):

I'll make sure that your kids hear about this in the future ~ ლ(́◉◞౪◟◉‵ლ)

OpenStudy (charlotte123):

AND YOUR DAM WIFE TOO ლ(́◉◞౪◟◉‵ლ)